OpenStudy (anonymous):

What is the slope and y-intersept of this line Y=-2x+8 a slope:8 y-intercept:-2 b slope : 2 y-intercept: -8 c slope: -1 y-intercept: 4 d slope: -2 y-intercept : 8

OpenStudy (anonymous):

Please help I have no clue how to do this

pooja195 (pooja195):

Ok your equation is written in "slope intercept form" y=mx+b = slope intercept form Where m is the slope B is the y intercept Can you identify the slope and y intercept using the info i provided above? Y=-2x+8

OpenStudy (anonymous):

8

pooja195 (pooja195):

What is 8 ? The y intercept or the slope?

OpenStudy (anonymous):

I think it is 8

pooja195 (pooja195):

Once again. What is 8 ? The y intercept or the slope?

OpenStudy (anonymous):

Y intercept I think not 100% sure

pooja195 (pooja195):

Correct 8 is the y intercept :) Whats the slope?

OpenStudy (anonymous):

-2

pooja195 (pooja195):

Correct! So which answer choice corresponds to what we decided? Slope=-2 y-intercept=8

OpenStudy (anonymous):

So D

pooja195 (pooja195):

Correct well done! :)

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (anonymous):

I have more questions can I just type them in here

OpenStudy (anonymous):

So D

OpenStudy (anonymous):

-2

pooja195 (pooja195):

Make a new post.

OpenStudy (anonymous):

I am trying to but my phone is messing up

pooja195 (pooja195):

Ok then post it here

OpenStudy (anonymous):

What is the slope and y intercept of this line y= x-3. A slope : -3 y intercept :1. B slope :1 y intercept :-3. C slope: 3 y intercept: -1. D slope: -1 y intercept: -3

pooja195 (pooja195):

Same concept but in the case the slope is x which means 1 so slope =1 figure out the y intercept :)

OpenStudy (anonymous):

-3

pooja195 (pooja195):

So the answer is?

OpenStudy (anonymous):

B

pooja195 (pooja195):

Good.

OpenStudy (anonymous):

Thank you for the help

OpenStudy (anonymous):

B

OpenStudy (anonymous):

I am trying to but my phone is messing up

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

I think it is 8

OpenStudy (anonymous):

Please help I have no clue how to do this

pooja195 (pooja195):

???

OpenStudy (anonymous):

I have no idea what happens here

pooja195 (pooja195):

oh >.>

OpenStudy (anonymous):

What is the slope - intercept form of this equation 8x+6y=12. A y=-4/3x+2. B y=-4/3x+6. C y=4/3x+6. D y=2/3x+2